# Towards a Theory of Pragmatic Information, with Applications

### v3

@@ 4 Pragmatic Utility and the Efficient Market Hypothesis
As indicated in the Introduction, a given message/message ensemble can, for a given decision maker, $\Delta$, have positive pragmatic information, yet misinform $\Delta$. This follows from our definition of pragmatic information as information that changes $\Delta$’s views of the state of the world, regardless of whether the new view is any more accurate, better, etc. than some base state. Furthermore, a knowledge of the decision facing $\Delta$ may not provide enough additional context to assign a value to that decision and thus to the messages leading to that decision. For example, what should be the value of a message that can prompt a single, terrible, but low probability outcome, as opposed to a message that leads to a set of more likely, but merely bad outcomes? Or can we be sure that we can assign values $V(m_{1}),V(m_{2}),V(m_{3})$ to messages $m_{1},m_{2},m_{3}\in\mathcal{M}$, such that $V(m_{1})<V(m_{2})$ and $V(m_{2})<V(m_{3})$ implies $V(m_{1})<V(m_{3})$?
@@
For a given $\Delta$, we can then partition $\mathcal{M}$ as $\mathcal{M}=\mathcal{I}_{\Delta}\cup\mathcal{H}_{\Delta}\cup\mathcal{U}_{\Delta}$, where:
@@
We now apply the above definitions to make a useful reformulation of the efficient market hypothesis, a cornerstone of financial economics (See, for example, [16].). According to this hypothesis, asset prices, such as stock and bond prices, reflect all available information, so there is nothing to be gained by poring over charts of price histories or over annual reports.
The standard quantitative statement of the efficient market hypothesis is, roughly speaking, that asset price fluctuations are Markov processes. However, the assumption that market efficiency is a property of asset price fluctuations suggests that all investors should find the markets equally inscrutable, a claim belied by the conspicuous long term success of such investors as Warren Buffett. If we take the pragmatic utility of market information to be its ability to improve its recipient’s ability to achieve above market returns, we can therefore propose the following reformulation:
@@
Clearly, the above characterization of market efficiency depends on the participant, as well as the message ensemble being considered. The market can therefore be efficient to some participants, but not others beause information may be pragmatically useful to one market participant but not others [^8] This is certainly the case, for example, regarding knowledge of short term mispricings that can be exploited by automated trading systems, but not human traders.. The market can also be efficient with respect to certain message ensembles, but not others. The notion of efficiency relative to a given dataset is already a staple of the relevant literature999[16], for example, distinguishes between the “weak”, “semi-strong”, and “strong” forms of the hypothesis (that, respectively, price histories, all available public information, and all available information, including insider information, are irrelevant in predicting future asset price fluctuations).. Nevertheless, our conceptual framework seems to provide one immediate insight: namely that a market is necessarily efficient with respect to a given participant and a given message ensemble if that participant is unable to extract any pragmatic information from that ensemble. One such situation is when the amount of incoming market information overwhelms the computational capacities of the participant, as the present author has found from personal experience!
@@ 5 Directions for Future Research: $\Delta$’s Computational Capacity and Beyond
The most compelling result of the Shannon theory, the so-called noisy coding theorem [13], does not seem to have an analogue for pragmatic information. In other words, it does not seem possible to formulate a “pragmatic channel capacity”, i.e. a rate above which pragmatic information cannot be transmitted. This is because a given message can have an arbitrarily small a priori probability of receipt, so that the logarithm in (2) can be made arbitrarily large with even a moderately likely a posteriori update.
Implicit in the above, however, are the limits that a decision maker’s computational abilities play in accruing pragmatic information. A sense of the importance of the role of these limits is made clear by an example due to [8]: If $\Delta$ has no computational abilities at all, in the sense that any message received will be ignored, $\mathbf{p}_{m}=\mathbf{q}$ for all messages $m$, and $\Phi_{\Delta}({\mathcal{M}};\Omega)=0$, regardless of the length or complexity of $m$! In other words, no computation implies no pragmatic information!
@@
All of the above suggests the need for a systematic exploration of how $\Delta$’s ability to accrue pragmatic information depends on its computational abilities. One approach might be to formalize these abilities via the Chomsky Hierarchy familiar to students of abstract automata [19] or the more detailed hierarchy in [18]. The latter author also suggests that the extensive semigroup theory of automata might be of use in understanding how $\Delta$ partitions its input messages into its possible updated “state of the world” estimates.

### v5

@@ 5 Pragmatic Utility and the Efficient Market Hypothesis
As indicated in the Introduction, a given message/message ensemble can, for a given decision maker, $\Delta$, have positive pragmatic information, yet misinform $\Delta$. This follows from our definition of pragmatic information as information that changes $\Delta$’s views of the state of the world, regardless of whether the new view is any more accurate, better, etc. than some base state. Nor is $\Delta$ guaranteed to have enough context to assign a value to that decision and thus to the messages leading to that decision. For example, what should be the value of a message that can prompt a single, terrible, but low probability outcome, as opposed to a message that leads to a set of more likely, but merely bad outcomes? Or can we be sure that we can assign values $V(m_{1}),V(m_{2}),V(m_{3})$ to messages $m_{1},m_{2},m_{3}\in\mathcal{M}$, such that $V(m_{1})<V(m_{2})$ and $V(m_{2})<V(m_{3})$ implies $V(m_{1})<V(m_{3})$?
@@
For a given $\Delta$, we can then partition $\mathcal{M}$ as $\mathcal{M}=\mathcal{I}_{\Delta}\cup\mathcal{D}_{\Delta}\cup\mathcal{U}_{\Delta}$, where:
@@
In particular, if the pragmatic information in some message ensemble $\mathcal{E}$ is zero, then it cannot contain any pragmatically useful information.
We now apply the above definitions to make a useful reformulation of the efficient market hypothesis, a cornerstone of financial economics (See, for example, [18].). According to this hypothesis, current asset prices, such as stock and bond prices, reflect all available information, so there is nothing to be gained by poring over charts of price histories or over annual reports.
However, the assumption that market efficiency is a property of asset price fluctuations suggests that all investors should find the markets equally inscrutable, a claim belied by the conspicuous long term success of such investors as Warren Buffett. If we take the pragmatic utility of market information to be its ability to improve its recipient’s ability to achieve above-market returns, we can therefore propose the following reformulation:
@@
Clearly, the above characterization of market efficiency depends on the participant, as well as the message ensemble being considered. The market can therefore be efficient to some participants, but not others because information may be pragmatically useful to one market participant but not others [^7] This is certainly the case, for example, regarding knowledge of short term mispricings that can be exploited by automated trading systems, but not human traders.. That the market can be efficient with respect to certain message ensembles, but not others is already a staple of the relevant literature. [18], for example, distinguishes between the “weak”, “semi-strong”, and “strong” forms of the hypothesis that, respectively, price histories, all available public information, and all available information, including insider information, are irrelevant in predicting future asset price fluctuations. Nevertheless, our conceptual framework seems to provide one immediate insight: namely that a market is necessarily efficient with respect to a given participant and a given message ensemble if that participant is unable to extract any pragmatically useful information from that ensemble. One such situation is when the amount of incoming market information overwhelms the computational capacities of the participant, as the present author has found from personal experience!
@@ 6 Directions for Future Research: $\Delta$’s Computational Capacity and Beyond
Implicit in the above are the limits that a decision maker’s computational abilities play in accruing pragmatic information. A sense of the importance of the role of these limits is made clear by an example due to [9]: If $\Delta$ has no computational abilities at all, in the sense that any message received will be ignored, $\mathbf{p}_{m}=\mathbf{q}$ for all messages $m$, and $\Phi_{\Delta}({\mathcal{M}};\Omega)=0$, regardless of the length or complexity of $m$! In other words, no computation implies no pragmatic information!
@@
All of the above suggests the need for a systematic exploration of how $\Delta$’s ability to accrue pragmatic information depends on its computational abilities. One approach might be to formalize these abilities via the Chomsky Hierarchy familiar to students of abstract automata [20] or the more detailed hierarchy in [1]. The latter author also suggests that the extensive semigroup theory of automata and his statistical mechanical approach might be of use in understanding how $\Delta$ partitions its input messages into its possible updated “state of the world” estimates.